A batter hits a baseball so that it leaves the bat with an initial velocity of 28.0m/s at an angle of 35°. What is the maximum height, the time it is in the air, and the range?

Single line text.



1
Expert's answer
2021-06-06T21:36:01-0400

The information given is

  • The initial velocity is "V_{o}=28.0\;\text{m}/\text{s}"
  • The launch angle is "\theta=35^{0}"
  • Gravity acceleration is "a_{y}=-9.81\;\text{m}/\text{s}^{2}"
  • The starting position for the move is "x_{o}=0\;\text{m}\;\text{and}\;y_{o}=0\;\text{m}"

\\

Components of the initial velocity.

Horizontal component.

"cos(35^{0})=\dfrac{V_{ox}}{V_{o}}\\
cos(35^{0})\;V_{o}=V_{0x}\\
V_{0x}=V_{o}\;cos(35^{0})\\
V_{ox}=28.0\;\text{m}/\text{s}\times cos(35^{0})\\
V_{ox}=22.9\;\text{m}/\text{s}"


Vertical component.

"sin(35^{0})=\dfrac{V_{oy}}{V_{o}}\\
sin(35^{0})\;V_{o}=V_{0y}\\
V_{0y}=V_{o}\;sin(35^{0})\\
V_{oy}=28.0\;\text{m}/\text{s}\times sin(35^{0})\\
V_{oy}=16.1\;\text{m}/\text{s}"

\\

The maximum height is calculated with the following equation.

"\Delta Y_{max}=\dfrac{ V_{fy}^{2}-V_{oy}^{2}}{2\;a_{y}}"

\\

Where.

"V_{oy}" is the initial vertical velocity.

"V_{fy}" is the velocity at the maximum height.

"a_{y}" is the acceleration of gravity.

\\

Remember that at the top, the vertical component of velocity is zero.

"V_{fy}=0\;\text{m}/\text{s}"

\\

Evaluating numerically.

"\Delta Y_{max}=\dfrac{ V_{fy}^{2}-V_{oy}^{2}}{2\;a_{y}}\\
\Delta Y_{max}=\dfrac{ (0\;\text{m}/\text{s} )^{2}-(16.1\;\text{m}/\text{s})^{2}}{2\times-9.81\;\text{m}/\text{s}^{2}}\\
\Delta Y_{max}=13.2\;\text{m}"

\\

!!Answer.

The maximum height is "\displaystyle \color{red}{\boxed{\Delta Y_{max}=13.2\;\text{m}}}"

\\

Flight time is calculated using the following equation.

"V_{fy}=V_{oy}+a_{y}\;t"

\\

where

"V_{oy}" is the initial vertical velocity.

"V_{fy}" is the final vertical velocity.

"a_{y}" is the vertical acceleration.

\\

Remember that if the movement is symmetrical, the vertical velocity with which it reaches the ground is of the same magnitude but in the opposite direction

"V_{fy}=-V_{oy}" .

\\

Obtaining the expression for time.

"V_{fy}=V_{oy}+a_{y}\;t\\
V_{oy}+a_{y}\;t=V_{fy}\\
a_{y}\;t=V_{fy}-V_{oy}\\
t=\dfrac{V_{fy}-V_{oy}}{a_{y} }\\
\text{if}\;V_{fy}=-V_{oy}\\
t=\dfrac{-V_{oy}-V_{oy}}{a_{y}}\\
t=\dfrac{-2\;V_{oy}}{a_{y}}"

\\

Evaluating numerically.

"t=\dfrac{-2\;V_{oy}}{a_{y}}\\
t=\dfrac{-2\times 16.1\;\text{m}/\text{s}}{-9.81\;\text{m}/\text{s}^{2}}\\
t=3.28\;\text{s}"

\\

!!Answer.

Flight time is "\displaystyle \color{red}{\boxed{t=3.28\;\text{s}}}"

\\

The distance traveled is given by the following equation.

"X=X_{0}+V_{ox}\;t"

\\

where.

"V_{ox}" is the horizontal velocity.

"t" Is time.

"X_{o}" is the starting position.

\\

Evaluating numerically

"X=X_{0}+V_{ox}\;t\\
X=0\;\text{m}+22.9\;\text{m}/\text{s}\times 3.28\;\text{s}\\
x=75.1\;\text{m}"

\\

!!Answer

The maximum horizontal reach is "\displaystyle \color{red}{\boxed{x=75.1\;\text{m}}}"